The nth term t(n),of a sequence is given by t(n)=p+nq,where p and q are constants.find the sum of the first 6 term in terms of p and q
Answer by greenestamps(13200)



t(1) = p+1q
t(2) = p+2q
t(3) = p+3q
t(4) = p+4q
t(5) = p+5q
t(6) = p+6q

Add those expressions by combining like terms to get the sum of t(1) through t(6) in terms of p and q.